# Deep reinforcement learning-based active flow control of an elliptical cylinder: transitioning from an elliptical cylinder to a circular cylinder and a flat plate

### v3

@@ III RESULTS
III.1 Baseline Flow
@@
Across all test cases, persistent instabilities in the form of vortices are observed at the exit of the computational domain, indicating that the fluid flow remains unstable in the downstream direction. As the fluid advances downstream, there is a tendency for the peak vortex intensity to decrease. This attenuation of vorticity is primarily attributed to the viscous diffusion effects commonly present in low Reynolds number regions. This finding is consistent with the studies by Johnson, Thompson, and Hourigan Johnson, Thompson, and Hourigan (2004). Between $Ar$ of 2 and 1.25, in addition to the recirculation bubble downstream of the cylinder, two additional recirculation bubbles form near the upstream wall. As $Ar$ is further reduced, the confined flow exhibits only one recirculation bubble. This observation is similar to that of Sahin and Owens Sahin and Owens (2004), where two additional recirculation bubbles appear on the solid sidewalls of the computational domain under the influence of blockage ratio and Reynolds number. These phenomena in the wake field of elliptical cylinders represent some intriguing physical behaviors. This serves as a foundation for subsequent flow control studies.
@@
III.2 Effectiveness of AFC
@@
For a detailed demonstration of vortex suppression, we present in Fig. 7 an example using an elliptical cylinder with $Ar=0.75$, observing the evolution of the instantaneous vorticity field throughout the DRL training process. We select four representative snapshots of the instantaneous vorticity field to depict changes in vortex shapes and positions. Labels A,B,C,D and E in Fig. 7(a) denote the five vortices initially shed. By the time depicted in Fig. 7(b), these initial five vortices have progressed downstream without new vortices being shed under the influence of flow control. In Fig. 7(c), labeled as A,B,C,D and E, the vortices have developed near the exit of the computational domain, exhibiting elongated shapes as they are about to be expelled from the domain. At this stage, the shape of the recirculation bubble has stabilized, and no new vortices are shed. By the time represented in Fig. 7(d), the originally labeled five vortices have all been expelled from the computational domain, the area of the recirculation bubble no longer expands, and the entire computational domain has reached a stable state. Under the conditions of $Ar=1.75$, 1.5, 1.25, and 1 for the elliptical cylinder, the DRL-based AFC demonstrates similar control effectiveness.
Figure 7: For an elliptical cylinder with $Ar=0.75$, observed the evolution of the instantaneous vorticity snapshots when DRL training. Specifically, focused on four representative training episodes: (a) Episode = 0, (b) Episode = 450, (c) Episode = 800, and (d) Episode = 2100. The initial shedding vortex are labeled as A, B, C, D and E, and them deformation and displacement are subsequently tracked. The variations in the instantaneous vorticity snapshots at these episodes describe the influence of flow control techniques on vortex shedding.
@@
Figure 10: Comparison of $C_{D}$ and $C_{L}$ before and after control. The background color represents the unstable stage of control, and the absence of background color means that the corresponding drag or lift has been in a stable stage during this period. (a) and (b) $Ar=2$; (c) and (d) $Ar=1.75$; (e) and (f) $Ar=1.5$; (g) and (h) $Ar=1.25$; (i) and (j) $Ar=1$; (k) and (l) $Ar=0.75$.
Further, Fig. 10 and Fig. 11 provide a comparative analysis of the $C_{D}$ and $C_{L}$ before and after implementing the DRL-based AFC strategy, showcasing its effectiveness across a range of $Ar$ from 2 to 0.75 and from 0.5 to 0, respectively. As shown in Fig. 10, the AFC strategy has no effect on the stability of both $C_{D}$ and $C_{L}$ at an $Ar$ of 2. However, when synthetic jets control is activated at $Ar$ values of 1.75, 1.5, 1.25, 1, and 0.75, there is an immediate and substantial reduction in $C_{D}$, which quickly reaches a minimum value. After half of the total control duration, $C_{D}$ stabilizes and remains constant at this minimum value. Similarly, $C_{L}$ experiences significant fluctuations initially but stabilizes near zero after half of the total control duration. The reductions in $C_{D}$ and $C_{L}$ at different $Ar$ values can be found in Tab. 2. Clearly, the periodic oscillations of $C_{D}$ and $C_{L}$ originally present in the baseline flow are significantly reduced, ultimately transitioning to a stable state.
@@
The DRL-based AFC implemented in our study is demonstrated across Figs. 8, 9, 10 and 11. This control strategy not only achieves high control performance but also maintains robust effectiveness, as evidenced by the quantified variations in $C_{D}$, $C_{L}$, and the ratio of synthetic jet mass flow rate to incoming flow rate, presented in Tab. 2. $\overline{C}_{D,\text{Baseline}}$ and $\overline{C}_{L,\text{Baseline}}$ represents the time-averaged $C_{D}$ and $C_{L}$ of the elliptical cylinder once the baseline flow has fully developed. $\overline{C}_{D,\text{Controlled}}$, $\overline{C}_{L,\text{Controlled}}$, and $\overline{a}$ denote the time-averaged values under stable conditions after AFC is applied, excluding the transient response phase. At $Ar=2$, the results of AFC control still maintain a stable baseline flow, without applying external excitation to disrupt flow stability. In the range of $Ar$ values from 1.75 to 0.75, the AFC control achieves substantial reductions in $C_{D}$ ranging from 0.9% to 15.7%, while the reduction rates in $C_{L}$ exceed 95% in all cases. Notably, the reduction in the coefficient of $C_{L}$ for elliptical cylinders, ranging from 99.7% to 95.2%, aligns with the achievement of a completely stable state (free from oscillations and fluctuations) for both $C_{D}$ and $C_{L}$ as illustrated in Fig. 10. This consistency is also mirrored in the suppression of alternating vortex shedding in the Baseline, as depicted in Fig. 8.
@@
The above content details the effectiveness of the flow control strategy, with the corresponding DRL training process curves for elliptical cylinders with different $Ar$ values presented in Fig. 13. First, we observed the overall trend of the reward function. For $Ar$ values ranging from 1.75 to 0, we noted that the reward function increases with the number of training iterations, eventually converging to a stable value. However, for $Ar=2$, after 5,000 training iterations, the reward function reaches a plateau but remains below zero, indicating that although the reward function has converged, the agent is unable to achieve the desired objective. Next, we examined the convergence performance of the reward function. For $Ar$ values ranging from 1.75 to 0, the reward function exhibits clear convergence points, stabilizing after a plateau phase. This indicates that the agent has learned effective strategies, achieving stable returns under these conditions. However, the convergence speed of DRL training varies significantly across different $Ar$ values. For $Ar$ values of 1.75, 1.5, and 1.25, convergence signs appear around 2,000 iterations. For $Ar$ values of 1 and 0.75, convergence signs emerge around 1,000 iterations. Despite setting the total training iterations to 5,000 for $Ar$ values between 1.75 and 0.75, we observed that the reward function converges much earlier for $Ar$ values of 1 and 0.75, indicating a faster convergence speed and the agent’s ability to quickly learn effective strategies. For $Ar=0.5$, convergence signs appear after 4,000 iterations. For $Ar=0.25$, convergence signs emerge after 8,000 iterations. For $Ar=0.1$, convergence signs are observed after 10,000 iterations. For $Ar=0$, convergence signs appear after 12,000 iterations. In the range of $Ar=0.5$ to 0, the DRL training convergence speed is significantly slower compared to $Ar$ values greater than 0.5. This reflects a lower training efficiency and greater difficulty in exploring effective strategies for the agent when $Ar$ is less than or equal to 0.5. Furthermore, in the case of different $Ar$, DRL training was conducted in parallel using 60 CFD environments, allowing the DRL algorithm to collect data simultaneously from multiple independent simulations. The parameters of the ANN were updated every 60 iterations, which means that the agent and the environment only interacted without updating the policy network for the first 60 episodes. After 60 episodes, the ANN network parameters started to update, resulting in an initial learning curve that exhibits a noticeable step-like increase. This parallel acceleration effect has also been observed in the research conducted by Rabault and KuhnleRabault and Kuhnle (2019a) and Wang and XuWang and Xu (2024a).
III.3 Analysis of the Controlled Flow
@@
Figure 14: When $Ar=2$, 1.75, 1.5, 1.25, 1, 0.75, the vortex shedding of the elliptical cylinder can be completely suppressed. The instantaneous vorticity snapshots of the wake flow field behind an elliptical cylinder, where the vortex shedding process is suppressed, feature key moments designated as $T_{1}$, $T_{2}$, and $T_{3}$. These represent the instances of the onset of control (when the vortex is fully developed), during the control process, and at the result of control, respectively.
@@
Figure 15: When $Ar=0.5$, 0.25, 0.1 and 0, vortex shedding from the elliptical cylinder is not entirely suppressed. The instantaneous vorticity snapshots of the wake flow field behind the elliptical cylinder capture key moments designated as $T_{1}$, $T_{2}$, and $T_{3}$. These correspond to the initiation of control (when the vortex is fully developed), the early phase of the control process, the mid-to-late phase of the control process, and the termination phase of control, respectively.

### v4

@@ III RESULTS
III.1 Flow characteristics of baseline flow
@@
Across all test cases, persistent instabilities in the form of vortices are observed at the exit of the computational domain, indicating that the fluid flow remains unstable in the downstream direction. As the fluid advances downstream, there is a tendency for the peak vortex intensity to decrease. This attenuation of vorticity is primarily attributed to the viscous diffusion effects commonly present in low Reynolds number regions. This finding is consistent with the studies by Johnson, Thompson, and Hourigan Johnson, Thompson, and Hourigan (2004). Between $Ar$ of 2 and 1.25, in addition to the recirculation bubble downstream of the cylinder, two additional recirculation bubbles form near the upstream wall. As $Ar$ is further reduced, the confined flow exhibits only one recirculation bubble. This observation is similar to that of Sahin and Owens Sahin and Owens (2004), where two additional recirculation bubbles appear on the solid sidewalls of the computational domain under the influence of blockage ratio and Reynolds number. These phenomena in the wake field of elliptical cylinders represent some intriguing physical behaviors.
This serves as a foundation for subsequent flow control studies.
@@
III.2 Training of DRL-based flow control
In this section, we use an elliptical cylinder with $Ar=0.75$ as an example to illustrate the interaction process between the agent and the environment during DRL training and the trial-and-error learning process to maximize the objective function. Fig. 7 shows the evolution of the vorticity field of the elliptical cylinder throughout the DRL training process, clearly demonstrating how the vortices shed from the elliptical cylinder in the baseline flow are gradually expelled from the computational domain while suppressing the emergence of new vortices during the execution of active flow control. During the training process, six characteristic episodes are marked, and the values of $C_{D}$, $C_{L}$, and the synthetic jet velocity for the elliptical cylinder are extracted for these episodes. Fig. 8 provides a more intuitive demonstration of the active flow control process and control strategies. Finally, Fig. 9 offers a detailed analysis of the streamline patterns before and after the execution of synthetic jet actions, with a local magnification of the streamlines near the synthetic jet location. This clearly shows the changes in streamlines around the elliptical cylinder when the synthetic jet is in the blowing or suction state. This section provides a comprehensive description of the DRL training process and the operational mechanism of the synthetic jet.
To facilitate understanding of how the agent learns the optimal flow control strategy through trial and error, we use the learning curve from the training process and the changes in the vorticity field around the elliptical cylinder to describe it. During DRL training, the agent first observes the state of the environment. We use several probes to provide physical information (velocity or pressure) of the flow field to characterize the state of the environment. The agent then calculates the initial reward function value based on the observed state and feeds the corresponding action (i.e., the jet velocity value) back to the CFD environment. Once the synthetic jets begin executing the action, they immediately influence the physical field around the elliptical cylinder, causing a change in its state. Next, the agent observes the state of the CFD environment again and calculates the reward value corresponding to that action. If the reward value increases, it indicates that the control strategy is beneficial for the objective function. If the reward value decreases, the agent will adjust the control strategy in the opposite direction. Through this iterative interaction and feedback process, the objective function is gradually optimized to the best control strategy.
Figure 7: For an elliptical cylinder with $Ar=0.75$, observed the evolution of the instantaneous vorticity snapshots when DRL training. Specifically, focused on four representative training episodes: (a) Episode = 1, (b) Episode = 650, (c) Episode = 1000, and (d) Episode = 2000. The variations in the instantaneous vorticity snapshots at these episodes describe the influence of flow control techniques on vortex shedding.
Figure 8: For an elliptical cylinder with $Ar=0.75$, the time series changes of the action, $C_{L}$, and $C_{D}$ during the characteristic episodes of DRL training. (a) The learning curve of DRL training for the elliptical cylinder with $Ar=0.75$. The values of the synthetic jet actions and the time series of $C_{L}$ and $C_{D}$ for episodes 1, 450, 650, 800, 1000, and 2000 are summarized in (b), (c), and (d), respectively. (b) The variation of the synthetic jet velocity with time. (c) The time series of $C_{L}$ for the characteristic episodes. (d) The time series of $C_{D}$ for the characteristic episodes. Note that the characteristic episodes in figures (b), (c), and (d) are consistent, with labels shown in (d).
Figure 9: Instantaneous streamlines of an elliptical cylinder with $Ar=0.75$. (a) In the baseline flow, streamline around the elliptical cylinder. To highlight the changes in streamlines during the synthetic jet blowing or suction process, localized streamline near 90° and 270° of the elliptical cylinder are presented on the right side. (b) In the baseline flow, with the synthetic jets turned off, the streamlines flow along the elliptical cylinder’s surface and separate near the aft side. (c) In the baseline flow, the fluid flows closely along the surface and separates near the aft side. (d) Instantaneous streamlines of the elliptical cylinder when flow control is activated. The synthetic jets perform suction on the upper side and blowing on the lower side of the elliptical cylinder. (e) and (f) display localized streamline near the synthetic jets. (e) Streamlines when the upper jet is performing suction. (f) Streamlines when the lower jet is performing blowing.
To demonstrate the suppression of wake vortices in an elliptical cylinder during DRL training, we provide an example using an elliptical cylinder with $Ar=0.75$ in Fig. 7, observing the evolution of the instantaneous vorticity field throughout the DRL training process. The DRL training consists of 6000 episodes, including a rapid growth phase and a convergence phase. After an initial rapid rise during the first 2000 episodes, the reward function value approaches its maximum. In the 2000-6000 episode phase, the reward function value fluctuates near its maximum, indicating that the training has stably converged. We select four typical episodes during the surge phase to describe the changes in vortex shapes and positions. Fig. 7(a) shows the wake field of the elliptical cylinder at episode 1, where regular, periodic vortex shedding is observed. DRL training begins with control applied to an already fully developed flow around the elliptical cylinder. In Fig. 7(b), no new vortices shed under the influence of synthetic jet blowing or suction, and the recirculation region behind the elliptical cylinder elongates. The initially unstable and oscillating recirculation region behind the elliptical cylinder becomes symmetric and stable during flow control execution. Fig. 7(c) shows that the vortices shed in the baseline flow gradually move downstream, stretching and interacting with each other. At this stage, the shape of the recirculation bubble stabilizes, and no new vortices shed. By episode 2000, Fig. 7(d) illustrates that no vortex shedding exists in the computational domain, the area of the recirculation bubble no longer expands, and the entire computational domain reaches a stable state. This indicates that when the DRL training reward value converges near its maximum, the agent has learned the optimal control strategy, effectively suppressing vortex shedding in the wake field through the actions of the synthetic jet.
To more clearly illustrate the variations in synthetic jet velocity, $C_{L}$, and $C_{D}$ during DRL training, we selected six characteristic episodes to extract the corresponding actions, $C_{L}$, and $C_{D}$. Fig. 8(a) divides the DRL training into two phases: the reward function value rapidly increases between episodes = 0 and episodes = 2000, and after episode = 2000, the reward function enters a plateau, indicating that the training has converged. We selected episodes = 1, episodes = 450, episodes = 650, episodes = 800, episodes = 1000, and episodes = 2000 as characteristic episodes to observe the changes in $C_{D}$, $C_{L}$, and the control strategies for the elliptical cylinder. Fig. 8(b) describes the variation of the synthetic jet velocity over time, representing the development of the flow control strategy. It can be observed that at episode 1, the agent attempts to use a relatively low jet velocity. As the number of episodes increases, the amplitude of the jet velocity gradually increases. By episode = 2000, the control strategy involves using a high jet velocity in the initial phase to control the flow field, followed by maintaining stability with much lower jet velocities. Fig. 8(c) shows the time series of $C_{L}$ for the elliptical cylinder. At episode = 1, $C_{L}$ exhibits periodic fluctuations. As the number of episodes increases, the amplitude of $C_{L}$’s periodic fluctuations decreases. By episode = 2000, $C_{L}$ shows significant fluctuations in the initial phase of jet activation, then quickly reduces to near zero, maintaining zero for more than half the time, indicating that the lift around the elliptical cylinder is fully controlled. Fig. 8(d) shows the time series of $C_{D}$ for the elliptical cylinder. At episode = 1, $C_{D}$ exhibits periodic oscillations. As the number of episodes increases, the amplitude and frequency of $C_{D}$ fluctuations decrease. By episode = 2000, $C_{D}$ only shows severe fluctuations at the beginning of control and then stabilizes near its minimum value.
In Fig. 9, we illustrate how the synthetic jets disrupt the streamlines along the sidewalls of an elliptical cylinder with $Ar=0.75$ during active flow control. Fig. 9(a) shows the streamlines around the elliptical cylinder without active flow control. Fig. 9(b) and Fig. 9(c) provide zoomed-in views of the streamlines near the elliptical cylinder at the locations of the synthetic jets, which are not active at this point. Without control, the streamlines follow the contours of the elliptical cylinder and separate near the aft side. Fig. 9(d) shows the streamlines around the elliptical cylinder with flow control activated. Fig. 9(e) and Fig. 9(f) again provide zoomed-in views of the streamlines at the synthetic jet locations, but this time with the jets active. When control is activated, the synthetic jets either blow or suction relative to the sidewalls, disrupting the previously continuous streamlines that followed the solid boundary of the elliptical cylinder. The noticable difference between uncontrolled baseline and controlled flow is that the flow stays attached after applying jet control. In Fig. 9(b) and Fig. 9(c), distinct flow separation can be observed in the rear-end. With synthetic jet, the flow stays attached after the suction/blowing point in Fig. 9(e) and Fig. 9(f). This illustrates the mechanism by which the synthetic jets influence the flow, breaking the streamlines and altering the flow pattern around the elliptical cylinder.
III.3 DRL-based AFC for various ellipses
@@
Figure 12: Comparison of $C_{D}$ and $C_{L}$ before and after control. The background color represents the unstable stage of control, while the absence of background color indicates that the corresponding $C_{D}$ or $C_{L}$ remains in a stable stage during this period. (a) and (b) $Ar=2$; (c) and (d) $Ar=1.75$; (e) and (f) $Ar=1.5$; (g) and (h) $Ar=1.25$; (i) and (j) $Ar=1$; (k) and (l) $Ar=0.75$.
Further, Fig. 12 and Fig. 13 provide a comparative analysis of the $C_{D}$ and $C_{L}$ before and after implementing the DRL-based AFC strategy, showcasing its effectiveness across a range of $Ar$ from 2 to 0.75 and from 0.5 to 0, respectively. The training duration $T_{\text{max}}$ is set to 2.5. Time is normalized using $T_{\text{max}}$, with the control period starting at Time = 0 and ending at Time = 1. As shown in Fig. 12, the AFC strategy has no effect on the stability of both $C_{D}$ and $C_{L}$ at an $Ar$ of 2. However, when synthetic jets control is activated at $Ar$ values of 1.75, 1.5, 1.25, 1, and 0.75, there is an immediate and substantial reduction in $C_{D}$, which quickly reaches a minimum value. After half of the total control duration, $C_{D}$ stabilizes and remains constant at this minimum value. Similarly, $C_{L}$ experiences significant fluctuations initially but stabilizes near zero after half of the total control duration. The reductions in $C_{D}$ and $C_{L}$ at different $Ar$ values can be found in Tab. 3. Clearly, the periodic oscillations of $C_{D}$ and $C_{L}$ originally present in the baseline flow are significantly reduced, ultimately transitioning to a stable state.
@@
The DRL-based AFC implemented in our study is demonstrated across Figs. 10, 11, 12 and 13. This control strategy not only achieves high control performance but also maintains robust effectiveness, as evidenced by the quantified variations in $C_{D}$, $C_{L}$, and the ratio of synthetic jet mass flow rate to incoming flow rate, presented in Tab. 3. $\overline{C}_{D,\text{Baseline}}$ and $\overline{C}_{L,\text{Baseline}}$ represents the time-averaged $C_{D}$ and $C_{L}$ of the elliptical cylinder once the baseline flow has fully developed. $\overline{C}_{D,\text{Controlled}}$, $\overline{C}_{L,\text{Controlled}}$, and $\overline{a}$ denote the time-averaged values under stable conditions after AFC is applied, excluding the transient response phase. At $Ar=2$, the results of AFC control still maintain a stable baseline flow, without applying external excitation to disrupt flow stability. In the range of $Ar$ values from 1.75 to 0.75, the AFC control achieves substantial reductions in $C_{D}$ ranging from 0.9% to 15.7%, while the reduction rates in $C_{L}$ exceed 95% in all cases. Notably, the reduction in the coefficient of $C_{L}$ for elliptical cylinders, ranging from 99.7% to 95.2%, aligns with the achievement of a completely stable state (free from oscillations and fluctuations) for both $C_{D}$ and $C_{L}$ as illustrated in Fig. 12. This consistency is also mirrored in the suppression of alternating vortex shedding in the Baseline, as depicted in Fig. 10.
@@
The above content details the effectiveness of the flow control strategy, with the corresponding DRL training process curves for elliptical cylinders with different $Ar$ values presented in Fig. 15. First, we observed the overall trend of the reward function. For $Ar$ values ranging from 1.75 to 0, we noted that the reward function increases with the number of training iterations, eventually converging to a stable value. However, for $Ar=2$, after 5,000 training iterations, the reward function reaches a plateau but remains below zero, indicating that although the reward function has converged, the agent is unable to achieve the desired objective. Next, we examined the convergence performance of the reward function. For $Ar$ values ranging from 1.75 to 0, the reward function exhibits clear convergence points, stabilizing after a plateau phase. This indicates that the agent has learned effective strategies, achieving stable returns under these conditions. However, the convergence speed of DRL training varies significantly across different $Ar$ values. For $Ar$ values of 1.75, 1.5, and 1.25, signs of convergence appear around 2,000 iterations. For $Ar$ values of 1 and 0.75, signs of convergence emerge around 1,000 iterations.
Despite setting the total training iterations to 5,000 for $Ar$ values between 1.75 and 0.75, we observed that the reward function converges much earlier for $Ar$ values of 1 and 0.75, indicating a faster convergence speed and the agent’s ability to quickly learn effective strategies. For $Ar=0.5$, convergence appears after 4,000 iterations. For $Ar=0.25$, convergence is reached after 8,000 iterations. For $Ar=0.1$, convergence can be observed after 10,000 iterations. For $Ar=0$, convergence appears after 12,000 iterations. In the range of $Ar=0.5$ to 0, the DRL training convergence speed is significantly slower compared to $Ar$ values greater than 0.5. This reflects a lower training efficiency and greater difficulty in exploring effective strategies for the agent when $Ar$ is less than or equal to 0.5. Furthermore, in the case of different $Ar$, DRL training was conducted in parallel using 60 CFD environments, allowing the DRL algorithm to collect data simultaneously from multiple independent simulations. The parameters of the ANN were updated every 60 iterations, which means that the agent and the environment only interacted without updating the policy network for the first 60 episodes. After 60 episodes, the ANN network parameters started to update, resulting in an initial learning curve that exhibits a noticeable step-like increase. This parallel acceleration effect has also been observed in the research conducted by Rabault and KuhnleRabault and Kuhnle (2019a) and Wang and XuWang and Xu (2024a).
III.4 Analysis of the Controlled Flow
@@
Figure 16: When $Ar=2$, 1.75, 1.5, 1.25, 1, 0.75, the vortex shedding of the elliptical cylinder can be completely suppressed. The instantaneous vorticity snapshots of the wake flow field behind an elliptical cylinder, where the vortex shedding process is suppressed, feature key moments designated as $T_{1}$, $T_{2}$, and $T_{3}$. These represent the instances of the onset of control (when the vortex is fully developed), during the control process, and at the result of control, respectively. (a) $Ar=2$, (b) $Ar=1.75$, (c) $Ar=1.5$, (d) $Ar=1.25$, (e) $Ar=1.0$, (f) $Ar=0.75$.
@@
Figure 17: When $Ar=0.5$, 0.25, 0.1 and 0, vortex shedding from the elliptical cylinder is not entirely suppressed. The instantaneous vorticity snapshots of the wake flow field behind the elliptical cylinder capture key moments designated as $T_{1}$, $T_{2}$, and $T_{3}$. These correspond to the initiation of control (when the vortex is fully developed), the early phase of the control process, the mid-to-late phase of the control process, and the termination phase of control, respectively. (a) $Ar=0.5$, (b) $Ar=0.25$, (c) $Ar=0.1$, (d) $Ar=0$.